OpenStudy (anonymous):

what is the solution of the proportion? x-9/3=10/19

Parth (parthkohli):

Hmm...why don't you cross multiply?

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {a \over b} = {x \over y} \Longrightarrow ay = bx }\)

OpenStudy (anonymous):

i dont know how to cross multiply with this type of fraction

Parth (parthkohli):

\( \color{Black}{\Rightarrow \large {x - 9 \over 3} = {10 \over 19} }\) Is this your question?

OpenStudy (anonymous):

If this is your question \[x- 9/3 =10/19\] then just set \[x= 10/19 + 9/3\] Then go on from there. But that really doesn't like like a proportion.

OpenStudy (anonymous):

\[19(x-9)=3*10\]